1. I need three measures of central tendency—the mean, the median, and the mode—are more appropriate for certain populations than others.
I need to look on the in ternet. For each type of measure, give two examples of populations where it would be the most appropriate indication of central tendency.
2. Find the mean, median, and mode of the following data set:
5 15 9 22 67 42 2 72 81 53 6 70 41 9 42 23
3. Sometimes, we can take a weighted approach to calculating the mean. Take our example of high temperatures in July. Suppose it was 98°F on 7 days, 96°F on 14 days, 88°F on 1 day, 100°F on 6 days and 102°F on 3 days. Rather than adding up 31 numbers, we can find the mean by doing the following:
Mean = ( 1 x 88 + 14 x 96 + 7 x 98 + 6 x 100 + 3 x 102) / 31
…where 1, 14, 7, 6, and 3 are the weights or frequency of a particular temperature’s occurrence. Then we divide by the total of number of occurrences.
Suppose we are tracking the number of home runs hit by the Boston Red Sox during the month of August:
Number of Games HRs Hit each Day
2 3
5 2
6 1
7 0
Using the weighted approach, calculate the average number of home runs per game hit by the Sox.

Answer by Edwin McCravy(20055)

1. I need three measures of central tendency—the mean, the median, and the mode—are more appropriate for certain populations than others.
I need to look on the internet. For each type of measure, give two examples of populations where it would be the most appropriate indication of central tendency.
`
1
Mean, the heights of male basketball players.
Median, the salaries of all employees in a corporation.
Mode, Hat or shoe sizes
`
2. Find the mean, median, and mode of the following data set:
5,15,9,22,67,42,2,72,81,53,6,70,41,9,42,23
`
To find the mean:
`
Add them:
5+15+9+22+67+42+2+72+81+53+6+70+41+9+42+23 = 559
`
Count the number of number. There are 16.
`
Mean = 559/16 = 34.9375
`
To find the median:
`
Arrange them in order from smallest to largest:
`
2,5,6,9,9,15,22,23,41,42,42,53,67,70,72,81
`
Count the number of numbers. There are 16.
Since there are an even number of numbers, we
average the middle two numbers, the 8th and the
9th ones, 23 and 41.
`
Median = (23+41)/2 = 64/2 = 32
`
To find the Mode:
`
Look for the most often occurring numbers.
`
9 occurs twice and 42 occurs twice. So since there
is a "tie", there are TWO modes, 9 and 42. The data
is said to be BIMODAL.
`
`
`
3. Sometimes, we can take a weighted approach to calculating the mean. Take our example of high temperatures in July. Suppose it was 98°F on 7 days, 96°F on 14 days, 88°F on 1 day, 100°F on 6 days and 102°F on 3 days. Rather than adding up 31 numbers, we can find the mean by doing the following:
Mean = ( 1 x 88 + 14 x 96 + 7 x 98 + 6 x 100 + 3 x 102) / 31
…where 1, 14, 7, 6, and 3 are the weights or frequency of a particular temperature’s occurrence. Then we divide by the total of number of occurrences.
Suppose we are tracking the number of home runs hit by the Boston Red Sox during the month of August:
Number of Games | HRs Hit each Day  Product 
       2        |      3               6
       5        |      2              10
       6        |      1               6 
       7        |      0               0
---------------------------------------------
        Totals         6              22

Using the weighted approach, calculate the average number
of home runs per game hit by the Sox

Mean = 22/6 = 3.6666666667

Edwin